OpenStudy (anonymous):

integral of xe^-10x dx

hartnn (hartnn):

use integration by parts.

OpenStudy (anonymous):

okay which would be u and dv

OpenStudy (turingtest):

almost always, if you have e^(f(x)) times x to some power n u=x^n dv=e^(f(x))